whitemonsterbunny17 (whitemonsterbunny17):

Can someone please check my answers? Thanks! (ノ^ヮ^)ノ*:・゚✧*:・゚✧

whitemonsterbunny17 (whitemonsterbunny17):

Write the following measurements in scientific notation: a. 22 000 mg 22 * 10^3 b. 0.005 km .5 * 10^3 c. 65 900 000 m 65.9 * 10^5 d. 0.000 003 7 kg 0.37 * 10^6

whitemonsterbunny17 (whitemonsterbunny17):

Did I get any right?

ganeshie8 (ganeshie8):

Nope

whitemonsterbunny17 (whitemonsterbunny17):

What?! D: What am I doing wrong?

ganeshie8 (ganeshie8):

scientific notation form : \[\large \times \cdot \times\times\times\times\times* 10^{\times\times\times\times}\]

ganeshie8 (ganeshie8):

there should be ONLY ONE digit before the decimal point

ganeshie8 (ganeshie8):

a. 22 000 mg 22 * 10^3 mg but this is NOT scientifc notation

OpenStudy (anonymous):

*and if I may add, everytime you write something in scientific notation, you should put a decimal point*

OpenStudy (anonymous):

b and d look carefully

myininaya (myininaya):

Here is two examples of numbers written in scientific notation: \[523=5.23 \times 10^2 \] \[0.525=5.25 \times 10^{-1}\]

myininaya (myininaya):

that one digit before the decimal point make sure to not choose it to be 0

whitemonsterbunny17 (whitemonsterbunny17):

Okay, that kind of makes sense.. So would a be 2.2 x 10^3 ?

ganeshie8 (ganeshie8):

yes, what about the units ? you don't want to change units to SI in scientific notation ? what does your textbook/teacher say ?

ganeshie8 (ganeshie8):

hey actually its a NO... 22 x 10^3 is not same as 2.2 x 10^3

whitemonsterbunny17 (whitemonsterbunny17):

It doesn't say anything about the SI units.

whitemonsterbunny17 (whitemonsterbunny17):

Is the 2.2 part right?

ganeshie8 (ganeshie8):

ok then less work for you :)

ganeshie8 (ganeshie8):

yes that part is right !

whitemonsterbunny17 (whitemonsterbunny17):

Oh good, so all I have wrong is how many places, then.. so wait, it should be to the fourth? So 2.2 x 10^4 ?

ganeshie8 (ganeshie8):

You got it !

whitemonsterbunny17 (whitemonsterbunny17):

YAY!!!! Finally!! XD

OpenStudy (anonymous):

*fireworks go off in the back*

ganeshie8 (ganeshie8):

next one can be tricky

whitemonsterbunny17 (whitemonsterbunny17):

So you count everything AFTER the decimal point as the number of digits?

OpenStudy (anonymous):

Yea

ganeshie8 (ganeshie8):

Exactly ! thats nice way to put it

whitemonsterbunny17 (whitemonsterbunny17):

Okay, cool. So for the next one... would it be 5 x 10 ^-3 ?

ganeshie8 (ganeshie8):

who ate the decimal point

whitemonsterbunny17 (whitemonsterbunny17):

pac man! XD jk

ganeshie8 (ganeshie8):

5.0 x 10^-3 is more like scientific form :)

whitemonsterbunny17 (whitemonsterbunny17):

oh, so you need to have a decimal point in the front all the time? (I hate decimals)

ganeshie8 (ganeshie8):

lol i was a pacman addict once

ganeshie8 (ganeshie8):

yes decimal point is a must for scientists

whitemonsterbunny17 (whitemonsterbunny17):

^ Same haha And okay

whitemonsterbunny17 (whitemonsterbunny17):

So 5.0 x 10^(-3)

ganeshie8 (ganeshie8):

looks good if we pretend we're not concerned with significant figures

ganeshie8 (ganeshie8):

significant figures is a more tricky business, save it for later :)

whitemonsterbunny17 (whitemonsterbunny17):

Significant figures? Does that deal with the measurements?

ganeshie8 (ganeshie8):

yes, but one thing at a time

ganeshie8 (ganeshie8):

you need to study scientific notation before touching significant figures topic

whitemonsterbunny17 (whitemonsterbunny17):

Well, there was a question before on this same study quiz that said something about it.

whitemonsterbunny17 (whitemonsterbunny17):

This thing isn't graded tho, it's just to be used as notes for the test.. I skipped that question, but is it something I should know?

ganeshie8 (ganeshie8):

Yes you should know, but I think we can work the rpesent problem without refering to them

whitemonsterbunny17 (whitemonsterbunny17):

Okay, cool. ^.^ Do you want me to post the others in separate threads? There's only c and d left for this question.

ganeshie8 (ganeshie8):

lets finish here

ganeshie8 (ganeshie8):

c. 65 900 000 m

ganeshie8 (ganeshie8):

you just need to figure out two things : 1) where to put the decimal point 2) exponent of 10

whitemonsterbunny17 (whitemonsterbunny17):

Umm... 65.9 x 10^6 ?

ganeshie8 (ganeshie8):

remember, only one digit to the left of decimal point

ganeshie8 (ganeshie8):

it should be 6.59 x 10^7 ?

OpenStudy (anonymous):

Hey remember to include the units

whitemonsterbunny17 (whitemonsterbunny17):

Ohhhhhhhh I get it!

whitemonsterbunny17 (whitemonsterbunny17):

So you can't have like: 22.19 x 10^4 Or 143.29 x 10^4 You can only have 1.294 x 10^4

ganeshie8 (ganeshie8):

thats right !!

whitemonsterbunny17 (whitemonsterbunny17):

Aha! Okay. So 6.59 x 10^7

ganeshie8 (ganeshie8):

this is just one of those times where you just need to follow the rules

ganeshie8 (ganeshie8):

yes!

whitemonsterbunny17 (whitemonsterbunny17):

Okay, so the last one is really confusing looking. Would it be 0.37 x 10^(-7) ?

ganeshie8 (ganeshie8):

first digit cannot be 0

ganeshie8 (ganeshie8):

so moce the decimial point to right by one more position

whitemonsterbunny17 (whitemonsterbunny17):

3.7 x 10^-7?

ganeshie8 (ganeshie8):

d. 0.000 003 7 kg is same as 0.00 003 7 x 10^(-1)kg is same as 0.0 003 7x 10^(-2) kg is same as 0.003 7 x 10^(-3)kg is same as 0.03 7 x 10^(-4)kg is same as 0.3 7 x 10^(-5)kg is same as 3. 7 x 10^(-6)kg

ganeshie8 (ganeshie8):

only the last line is in scientific form

whitemonsterbunny17 (whitemonsterbunny17):

Okay, so for the powers, you only count the number of zeroes?

ganeshie8 (ganeshie8):

n * 10^3 : shifts the decimal point of number n to RIGHT by 3 positions n * 10^(-3) : shifts the decimal point of number n to LEFT by 3 positions

OpenStudy (anonymous):

i will leave this to @ganeshie8, must better idea lol

ganeshie8 (ganeshie8):

im bit dyslexic about left and right, so i always need to double check 2-3 times when messing with powers

whitemonsterbunny17 (whitemonsterbunny17):

Ohhh okay.

ganeshie8 (ganeshie8):

xx.xxxx * 10^3 = xxxxx.x

ganeshie8 (ganeshie8):

xx.xxxx * 10^(-3) = 0.0xxxxxx

whitemonsterbunny17 (whitemonsterbunny17):

Okay, that makes sense. I gotta go for a bit, supper, but I'll be back. ^.^ \(\huge\bf\rlap{\color{red}{\bigstar~T}\color{gold}{H}\color{greenyellow}{A}\color{lime}{N}\color{cyan}{K~}\color{blue}{Y}\hspace{0 pt}\color{purple}{O}\color{magenta}{U}\color{orchid}{!~\bigstar}}{\color{black}{\; \bigstar~THANK~YOU!~\bigstar}}\) Is it okay if I tag you in some other questions when I get back? (:

ganeshie8 (ganeshie8):

yeah sure :)

whitemonsterbunny17 (whitemonsterbunny17):

Awesome!! Thanks! :D

OpenStudy (anonymous):

And remember! I'm still here and I can help! :D

OpenStudy (anonymous):

*this is one of those awkward moments when i'm a third wheel XD*

ganeshie8 (ganeshie8):

3 wheels is more stable than 2 wheels :)